OpenStudy (anonymous):

What is the derivative of y=(x^2-1)^3

OpenStudy (accessdenied):

Do you know how to use the chain rule? y = f(g(x)) y' = f'(g(x)) * g'(x)

OpenStudy (anonymous):

somewhat

OpenStudy (rogue):

\[\frac {d}{dx} u^n = n u^{n-1} \frac {du}{dx}\]\[\frac {d}{dx} \left[ y=(x^2-1)^3 \right]\]\[\frac {dy}{dx} = 3(x^2 - 1)^2 \times \frac {d}{dx} (x^2 - 1)\]\[\frac {dy}{dx} = 6x(x^2 - 1)^2\]

OpenStudy (accessdenied):

Yes, that would be correct for the derivative. :P If we look at the f(x) to be x^3 and the g(x) to be (x^2 - 1), we'd have f(g(x)) = f(x^2 - 1) = (x^2 - 1)^3 The derivative of f(x) is 3x^2 f'(x) = 3x^2 f'(g(x)) = 3(x^2 - 1)^2 The derivative of g(x) would be 2x If we use that "y' = f'(g(x)) * g'(x)", we'd just insert those back in... y' = 3(x^2 - 1)^2 * (2x) = 6x(x^2 - 1)^2